# Synchronization of Dirac-Bianconi driven oscillators

### v1

@@ III.1 Example of a Dirac-Bianconi driven oscillator
As an example, let us consider the following Dirac-Bianconi system on a $1$-dimensional simplicial complex, i.e., a network, of $N_{n}$ 00-simplices, i.e., the nodes, and $N_{l}$ $1$-simplices, i.e., the links. We denote the 00-cochains with the variables $u_{i}$, for $i=1,..,N_{n}$, and the $1$-cochains with $v_{j}$, for $j=1,...,N_{l}$. The equations are given by the following
@@
More specifically, the model we choose is inspired by the celebrated FitzHugh-Nagumo model fitzhugh ; nagumo , a paradigmatic model in the study of neuronal dynamics Murray2001 . Such model consists of two variables: the first is a fast variable (in the definition of slow-fast dynamics kuehn2015multiple ) describing the neuron’s action potential through a cubic function, while the second is a slow variable modeling the ion density through a linear function. For certain values of the parameters, the FitzHugh-Nagaumo model exhibits periodic oscillations and can be analyzed through phase reduction, both as a single oscillator and as a network nakao16 ; nakao2018phase . In our setting, we consider the fast variables as 00-cochains, i.e., state variables on the nodes, while the slow variables as $1$-cochains, i.e., on the links. From a modeling point of view, this could make sense because we are considering two variables, the action potential and the recovery variable (related to the ion density), which have an intrinsically different nature and, hence, can be modeled as defined on different manifolds. Moreover, one can consider each neuron as a node with its action potential and each link as an axon or a dendrite with its ion concentration, which affects adjacent nodes. Note that this is not yet a realistic setting, given, among other things, that axons and dendrites are directed. The dynamical equations of the Dirac-Bianconi version of the FitzHugh-Nagumo model, which, for the sake of brevity, we will call DBFHN from now on, are given by the following
@@
Figure 2: Dynamics of the Dirac-Bianconi driven oscillator described in Eq. (9): in panel a), we depict the time series of the 00- and $1$-cochains; after a transient, the system follows a periodic trajectory of period $T=21.14$ time units (t.u.). Panel b) shows a snapshot of the values of the signals on nodes and links for $t=290$ t.u. The parameters are the following: $a_{1}=a_{2}=0.7$, $a_{3}=-3$, $b_{1}=b_{2}=b_{3}=0.3$, $\delta_{1}=\delta_{2}=\delta_{3}=0.08$, $I_{1}=I_{2}=I_{3}=0.8$. Note that $b_{j}$, $\delta_{j}$ and $I_{i}$ do not have to be equal for every system to have a periodic solution.
@@
Figure 3: Dynamics of each unit of the Dirac-Bianconi driven oscillator over one period: $\bm{B}$ is such that the system has a homogeneous state and the model parameters are $a_{1}=a_{2}=0.7$, $a_{3}=-3$, $b_{1}=b_{2}=b_{3}=0.3$, $\delta_{1}=\delta_{2}=\delta_{3}=0.08$, $I_{1}=I_{2}=I_{3}=0.8$ (same setting of the previous Figure). The resulting period of the Dirac-Bianconi driven oscillator is $T=21.14$ t.u.
In Fig. 3, we show the time series of each variable of the DBFHN of Fig. 2 over one period. Let us remark that, although each variable has a periodic behavior, the limit cycle lies in the $N_{n}+N_{l}$ phase space. In fact, any cochain would not oscillate without the (Dirac-Bianconi) interaction with the others, and it is the whole system that behaves as an oscillator and not any single 00- and $1$-cochain. Note that this approach, which has been considered on networks where oscillations arise only due to the interactions between the units wilson1972excitatory ; brunel1999fast ; kawamura2011collective ; pazo2016quasiperiodic ; devalle2017firing ; nakao2018phase ; politi2018collective ; di2018transition ; nakao2021sparse , is radically different from what generally found in the literature on synchronization, where each unit (i.e., the system on the node) is an oscillator by itself and one studies how the interactions between many oscillators affect the global dynamics pikovsky2001synchronization ; arenas2008synchronization .
As previously stated, this is a very simple setting: in fact, we have a $1$-simplicial complex of small size, the boundary operator with the orientation admitting a homogeneous state, only two state variables (one on the nodes and one on the links) and only one nonlinearity (for the dynamics of the 00-cochain). One could extend this set up in several ways, for example, by considering nonlinear functions for every $s$-cochain, which could also be highly dimensional, more than $2$ kinds of cochains (e.g., on nodes, links and triangles), a more complex structure of the simplicial complex, or even cell-complexes, which are obtained with regular politopes instead of simplices carletti2023global , etc. Nonetheless, the substance of what we will present in the following Sections would not change. The simplicity of this setting in which collective oscillations are observed is aimed at highlighting the framework and the approach, which are the main contributions of this work.
@@ III.2 Coupled Dirac-Bianconi driven oscillators
$\begin{cases}\dot{\bar{u}}_{i}=\bar{u}_{i}-\bar{u}_{i}^{3}-(\bar{\bm{B}}_{1}\vec{\bar{v}})_{i}+\bar{I}_{i},\\ \dot{\bar{v}}_{j}=\bar{\delta}_{j}((\bar{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j}-\bar{b}_{j}\bar{v}_{j}+\bar{\alpha}_{j}\bar{a}_{j}),\end{cases}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{ and }\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \begin{cases}\dot{\tilde{u}}_{i}=\tilde{u}_{i}-\tilde{u}_{i}^{3}-(\tilde{\bm{B}}_{1}\vec{\tilde{v}})_{i}+\tilde{I}_{i},\\ \dot{\tilde{v}}_{j}=\tilde{\delta}_{j}((\tilde{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j}-\tilde{b}_{j}\tilde{v}_{j}+\tilde{\alpha}_{j}\tilde{a}_{j}).\end{cases}$ (10)
@@
$\bar{\bm{B}}_{1}=\tilde{\bm{B}}_{1}=\begin{bmatrix}1&0&-1\\ -1&1&0\\ 0&-1&1\end{bmatrix},$ (11)
and $\bar{\alpha}_{j},\tilde{\alpha}_{j}=1$ from now on. Both systems have the same parameters as in the previous section with the only difference that $\tilde{I}_{3}=0.6$. With such parameters and without coupling, the period of DB1 is $\bar{T}=21.140$ time units (t.u.), while that for DB2 is $\tilde{T}=20.415$ t.u., meaning that their oscillation frequencies are $\bar{\omega}=0.2972$ t.u.-1 and $\tilde{\omega}=0.3078$ t.u.-1, respectively (hence, with a difference of $0.0906$ t.u.-1).
Figure 4: Weak interaction between two Dirac-Bianconi driven oscillators: pictorial representation of the setting considered in the text, namely, two Dirac-Bianconi driven oscillators weakly coupled via a pair of nodes. The pairwise coupling is, moreover, considered as linear and diffusive-like. Of course, more complex coupling configurations can be considered. Although it may seem that no higher-order effects take place in the coupling, synchronization between the two oscillators can be achieved only when considering the effects of the $1$-cochains, i.e., the variables on the links.
@@
$\begin{cases}\displaystyle\dot{\bar{u}}_{i}=\bar{u}_{i}-\bar{u}_{i}^{3}-(\bar{\bm{B}}_{1}\vec{\bar{v}})_{i}+\bar{I}_{i}+\epsilon\bar{g}_{i}(\bar{u}_{i},\tilde{u}_{i_{1}},...,\tilde{u}_{i_{N_{n}}}),\\ \dot{\bar{v}}_{j}=\bar{\delta}_{j}((\bar{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j}-\bar{b}_{j}\bar{v}_{j}+\bar{a}_{j}),\end{cases}$ (12)
$\begin{cases}\displaystyle\dot{\tilde{u}}_{i}=\tilde{u}_{i}-\tilde{u}_{i}^{3}-(\tilde{\bm{B}}_{1}\vec{\tilde{v}})_{i}+\tilde{I}_{i}+\epsilon\tilde{g}_{i}(\tilde{u}_{i},\bar{u}_{i_{1}},...,\bar{u}_{i_{N_{n}}}),\\ \dot{\tilde{v}}_{j}=\tilde{\delta}_{j}((\tilde{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j}-\tilde{b}_{j}\tilde{v}_{j}+\tilde{a}_{j}),\end{cases}$ (13)
@@
where $\bar{g}_{i},\tilde{g}_{i}$ are the coupling functions, and $\epsilon$ is the coupling strength, We have chosen the coupling through nodes as the simplest and, possibly, physically feasible way to couple Dirac-Bianconi driven oscillators. In fact, we consider each Dirac-Bianconi driven oscillator as an independent dynamical unit exhibiting complex (oscillatory) dynamics induced by the Dirac-Bianconi coupling, while the coupling between the oscillators is modeled with a classic diffusive coupling via the variables on the nodes (00-cochains). In principle, we could consider more elaborated higher-order couplings, however a pairwise weak ( $\leavevmode\nobreak\ 10^{-2}$) coupling allows us to focus on the dynamics of the Dirac-Bianconi driven oscillators. We assume the functions $\bar{g}_{i}$ and $\tilde{g}_{i}$ to be linear and diffusive. For the setting depicted in Fig. 4, where the coupled variables are $\bar{u}_{1}$ and $\tilde{u}_{1}$, the only non-zero coupling functions are $\bar{g}_{i}=\tilde{u}_{1}-\bar{u}_{1}=-\tilde{g}_{i}$. Of course, different coupling configurations could be considered, as well as different kinds of pairwise couplings, including nonlinear and non-diffusive ones. Again, we choose a simple setting in order to illustrate the phenomenon and the method.
@@
As discussed previously, the two DBFHN oscillators have slightly different frequencies, meaning that an incoherent behavior is obtained without coupling, i.e., $\epsilon=0$. If we now set $\epsilon=10^{-2}$ and we consider Eqs. (12) and (13) coupled via only one node (as depicted in Fig. 4), we achieve synchronization only for large values of the coupling strength, as shown in Fig. 5, where we observe that the two oscillators synchronize their frequencies as the coupling becomes stronger. An analogous behavior is observed for other linear diffusive coupling combinations of different pairs of nodes (results not shown).
@@
Figure 5: Synchronization of two Dirac-Bianconi driven oscillators weakly diffusively coupled: we show the frequencies of DB1 (blue) and DB2 (red), coupled as in Eqs. (12) and (13) integrated until $t=6000$ t.u. and averaged over $100$ periods. We can observe that, with diffusive coupling, synchronization is achieved only for large values of the coupling strength. The parameters are $\bar{a}_{1}=\bar{a}_{2}=\tilde{a}_{1}=\tilde{a}_{2}=0.7$, $\bar{a}_{3}=\tilde{a}_{3}=-3$, $\bar{b}_{1}=\bar{b}_{2}=\bar{b}_{3}=\tilde{b}_{1}=\tilde{b}_{2}=\tilde{b}_{3}=0.3$, $\bar{\delta}_{1}=\bar{\delta}_{2}=\bar{\delta}_{3}=\tilde{\delta}_{1}=\tilde{\delta}_{2}=\tilde{\delta}_{3}=0.08$, $\bar{I}_{1}=\bar{I}_{2}=\bar{I}_{3}=\tilde{I}_{1}=\tilde{I}_{2}=0.8$ and $\tilde{I}_{3}=0.4$.
@@
In addition to the diffusive coupling, let us now consider also a Dirac-Bianconi coupling. We can impose that the diffusive-like coupling between 00-cochains has an effect on the $1$-cochains lying on the links (i.e., $1$-chains) adjacent to the coupled nodes (i.e., 00-chains); vice versa, the $1$-cochains have an effect on the 00-cochains lying on the nodes. If we look again at Fig. 4, this means that, via a Dirac-Bianconi coupling, $\bar{u}_{1}$ affects $\tilde{v}_{1}$ and $\tilde{v}_{3}$, and $\tilde{u}_{1}$ affects $\bar{v}_{1}$ and $\bar{v}_{3}$; vice versa $\tilde{v}_{1}$ and $\tilde{v}_{3}$ affect $\bar{u}_{1}$, and $\bar{v}_{1}$ and $\bar{v}_{3}$ affect $\tilde{u}_{1}$. We represent the higher-order coupling functions with $h^{\mathcal{D}}$.
@@
$\begin{cases}\displaystyle\dot{\bar{u}}_{i}=\bar{u}_{i}-\bar{u}_{i}^{3}-(\bar{\bm{B}}_{1}\vec{\bar{v}})_{i}+\bar{I}_{i}+\epsilon\bar{g}_{i}(\bar{u}_{i},\tilde{u}_{i_{1}},...,\tilde{u}_{i_{N_{n}}})+\epsilon\bar{h}^{\mathcal{D}}_{i}\Big{(}\bar{u}_{i},(\bar{\bm{B}}_{1}\vec{\tilde{v}})_{i_{1}},...,(\bar{\bm{B}}_{1}\vec{\tilde{v}})_{i_{N_{n}}}\Big{)},\\ \\ \displaystyle\dot{\bar{v}}_{j}=\bar{\delta}_{j}((\bar{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j}-\bar{b}_{j}\bar{v}_{j}+\bar{a}_{j})+\epsilon\bar{h}^{\mathcal{D}}_{j}\Big{(}\bar{v}_{j},(\bar{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j_{1}},...,(\bar{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j_{N_{l}}}\Big{)},\end{cases}$ (14)
@@
$\begin{cases}\displaystyle\dot{\tilde{u}}_{i}=\tilde{u}_{i}-\tilde{u}_{i}^{3}-(\tilde{\bm{B}}_{1}\vec{\tilde{v}})_{i}+\tilde{I}_{i}+\epsilon\tilde{g}_{i}(\tilde{u}_{i},\bar{u}_{i_{1}},...,\bar{u}_{i_{N_{n}}})+\epsilon\tilde{h}^{\mathcal{D}}_{i}\Big{(}\tilde{u}_{i},(\tilde{\bm{B}}_{1}\vec{\bar{v}})_{i_{1}},...,(\tilde{\bm{B}}_{1}\vec{\bar{v}})_{i_{N_{n}}}\Big{)},\\ \\ \displaystyle\dot{\tilde{v}}_{j}=\tilde{\delta}_{j}((\tilde{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j}-\tilde{b}_{j}\tilde{v}_{j}+\tilde{a}_{j})+\epsilon\tilde{h}^{\mathcal{D}}_{j}\Big{(}\tilde{v}_{j},(\tilde{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j_{1}},...(\tilde{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j_{N_{l}}}\Big{)}.\end{cases}$ (15)
@@
In this case, the system achieves synchronization even for weak coupling strengths and, moreover, as the coupling strength grows, the oscillations are slowed down, as shown in Fig. 6. Hence, in our setting the Dirac-Bianconi coupling  [^5] Let us observe that the Dirac-Bianconi coupling of the previous equation would also emerge in the phase reduced model (see next Section) if the second order phase reduction leon19 of Eqs. (12) and (13) is performed, but it would be of order $\epsilon^{2}$. is fundamental in achieving synchronization for weak couplings. In the next section we will explain this behavior through the theory of phase reduction.
@@
Figure 6: Synchronization of two Dirac-Bianconi driven oscillators weakly coupled via the Dirac-Bianconi operator: we show the frequencies of DB1 (blue) and DB2 (red), coupled as in Eqs. (14) and (15) integrated until $t=6000$ t.u. and averaged over $100$ periods. On the left panel, we can observe that synchronization is achieved for small values of the coupling. The right panel shows that the Dirac-Bianconi coupling slows down the oscillations as the coupling strength is increased. Note the different scales of the $x$ axis in the two panels. The parameters are $\bar{a}_{1}=\bar{a}_{2}=\tilde{a}_{1}=\tilde{a}_{2}=0.7$, $\bar{a}_{3}=\tilde{a}_{3}=-3$, $\bar{b}_{1}=\bar{b}_{2}=\bar{b}_{3}=\tilde{b}_{1}=\tilde{b}_{2}=\tilde{b}_{3}=0.3$, $\bar{\delta}_{1}=\bar{\delta}_{2}=\bar{\delta}_{3}=\tilde{\delta}_{1}=\tilde{\delta}_{2}=\tilde{\delta}_{3}=0.08$, $\bar{I}_{1}=\bar{I}_{2}=\bar{I}_{3}=\tilde{I}_{1}=\tilde{I}_{2}=0.8$ and $\tilde{I}_{3}=0.4$.

### v2

@@ III.1 Example of a Dirac-Bianconi driven oscillator
As an example, let us consider the following Dirac-Bianconi system on a $1$-dimensional simplicial complex, i.e., a network, of $N_{n}$ 0-simplices, i.e., the nodes, and $N_{l}$ 11-simplices, i.e., the links. We denote the 0-cochains with the variables $u_{i}$, for $i=1,..,N_{n}$, and the $1$-cochains with $v_{j}$, for $j=1,...,N_{l}$. The equations are given by the following
@@
More specifically, the model we choose is inspired by the celebrated FitzHugh-Nagumo (FHN) model FitzHugh (1961); Nagumo et al. (1962), a paradigmatic model in the study of neuronal dynamics Murray (2001). Such model consists of two variables: the first is a fast variable (in the definition of slow-fast dynamics Kuehn (2015)) describing the neuron’s action potential through a cubic function, while the second is a slow variable modeling the ion density through a linear function [^5] Note that whether the FHN behaves as a fast-slow system depends on the time scale separation parameter, which we will denote with $\delta$ in what follows.. Fast-slow systems are widely studied in mathematical neuroscience because the time scales of the firing of the neuron and the changes in the ion channels are very different. Moreover, such models exhibit rich dynamics, such as canard cycles and Mixed Modes Oscillations (MMOs), which reproduce well the spiking of neurons Desroches et al. (2012, 2016). The FHN model can also exhibit periodic oscillations for certain parameter values and can be analyzed through phase reduction, both as a single oscillator and as a network Nakao (2016); Nakao et al. (2018); Zhu et al. (2020). In our setting, we consider the fast variables as 0-cochains, i.e., state variables on the nodes, while the slow variables as $1$-cochains, i.e., on the links. Note that, while considering the 0-cochains as the voltage of the neuron, the $1$-cochains do not find a direct correspondence with the recovery variable of the ion density of the node-based FHN model, for instance, because they are not segregated in the links. Indeed, this is not yet a realistic setting, given that, among other things, axons and dendrites are not symmetric links, but have a directionality. Nonetheless, we can think of $1$-cochains as a gradient slow variable, which triggers the spiking of the action potential; moreover, we believe that this can the starting point for modeling neuronal dynamics through this framework. The dynamical equations of the Dirac-Bianconi version of the FHN model, which, for the sake of brevity, we will call DBFHN from now on, are given by the following equations:
@@
Figure 2: Dynamics of the Dirac-Bianconi driven oscillator described in Eq. (9): in panel a), we depict the time series of the 0- and 11-cochains; after a transient, the system follows a periodic trajectory of period $T=21.14$ time units (t.u.). Panel b) shows a snapshot of the values of the signals on nodes and links for $t=290$ t.u. The parameters are the following: $a_{1}=a_{2}=0.7$, $a_{3}=-3$, $b_{1}=b_{2}=b_{3}=0.3$, $\delta_{1}=\delta_{2}=\delta_{3}=0.08$, $I_{1}=I_{2}=I_{3}=0.8$. Note that $b_{j}$, $\delta_{j}$ and $I_{i}$ do not have to be equal for every system to have a periodic solution.
@@
Figure 3: Dynamics of each unit of the Dirac-Bianconi driven oscillator over one period: $\bm{B}$ is such that the system has a homogeneous state and the model parameters are $a_{1}=a_{3}=0.7$, $a_{2}=-3$, $b_{1}=b_{2}=b_{3}=0.3$, $\delta_{1}=\delta_{2}=\delta_{3}=0.08$, $I_{1}=I_{2}=I_{3}=0.8$ (same setting of the previous Figure). The resulting period of the Dirac-Bianconi driven oscillator is $T=21.14$ t.u.
In Fig. 3, we show the time series of each variable of the DBFHN of Fig. 2 over one period. Let us remark that, although each variable has a periodic behavior, the limit cycle lies in the $N_{n}+N_{l}$ phase space. In fact, any cochain would not oscillate without the (Dirac-Bianconi) interaction with the others, and it is the whole system that behaves as an oscillator and not any single 0- and $1$-cochain. Note that this approach, which has been considered on networks where oscillations arise only due to the interactions between the units Wilson and Cowan (1972); Brunel and Hakim (1999); Kawamura et al. (2011); Pazó and Montbrió (2016); Devalle et al. (2017); Nakao et al. (2018); Politi et al. (2018); Di Volo and Torcini (2018); Nakao et al. (2021), is radically different from what generally found in the literature on synchronization, where each unit (i.e., the system on the node) is an oscillator by itself and one studies how the interactions between many oscillators affect the global dynamics Pikovsky et al. (2001); Arenas et al. (2008).
As previously stated, this is a very simple setting: in fact, we have a $1$-simplicial complex of small size, the boundary operator with the orientation admitting a homogeneous state, only two state variables (one on the nodes and one on the links) and only one nonlinearity (for the dynamics of the 0-cochain). One could extend this set up in several ways, for example, by considering nonlinear functions for every $s$-cochain, which could also be highly dimensional, more than $2$ kinds of cochains (e.g., on nodes, links and triangles), a more complex structure of the simplicial complex, or even cell-complexes, which are obtained with regular politopes instead of simplices Carletti et al. (2023), etc. Nonetheless, the substance of what we will present in the following Sections would not change. The simplicity of this setting in which collective oscillations are observed is aimed at highlighting the framework and the approach, which are the main contributions of this work.
@@ III.2 Coupled Dirac-Bianconi driven oscillators
[…]
@@
$\bar{\bm{B}}_{1}=\tilde{\bm{B}}_{1}=\begin{bmatrix}1&0&-1\\ -1&1&0\\ 0&-1&1\end{bmatrix}$. (11)
Both systems have the same parameters as in the previous section with the only difference that $\tilde{I}_{3}=0.6$. With such parameters and without coupling, the period of DB1 is $T_{1}$=21.140=21.140 time units (t.u.), while that for DB2 is $T_{2}$=20.415=20.415 t.u., meaning that their oscillation frequencies are $\omega_{1}$=0.2972=0.2972 t.u.-1 and $\omega_{2}$=0.3078=0.3078 t.u.-1, respectively (hence, with a difference of $0.0106$ t.u.-1).
Figure 4: Weak interaction between two Dirac-Bianconi driven oscillators: pictorial representation of the setting considered in the text, namely, two Dirac-Bianconi driven oscillators weakly coupled via a pair of nodes. The pairwise coupling is, moreover, considered as linear and diffusive-like. Of course, more complex coupling configurations can be considered. Although it may seem that no higher-order effects take place in the coupling, synchronization between the two oscillators can be achieved only when considering the effects of the 11-cochains, i.e., the variables on the links.
@@
$\begin{cases}\displaystyle\dot{\bar{u}}_{i}=\bar{u}_{i}-\bar{u}_{i}^{3}-(\bar{\bm{B}}_{1}\vec{\bar{v}})_{i}+\bar{I}_{i}+\epsilon\bar{g}_{i}(\bar{u}_{i},\tilde{u}_{i_{1}},...,\tilde{u}_{i_{N_{n}}})\\ \dot{\bar{v}}_{j}=\bar{\delta}_{j}((\bar{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j}-\bar{b}_{j}\bar{v}_{j}+\bar{\alpha}_{j}\bar{a}_{j})\end{cases},$ (12)
$\begin{cases}\displaystyle\dot{\tilde{u}}_{i}=\tilde{u}_{i}-\tilde{u}_{i}^{3}-(\tilde{\bm{B}}_{1}\vec{\tilde{v}})_{i}+\tilde{I}_{i}+\epsilon\tilde{g}_{i}(\tilde{u}_{i},\bar{u}_{i_{1}},...,\bar{u}_{i_{N_{n}}})\\ \dot{\tilde{v}}_{j}=\tilde{\delta}_{j}((\tilde{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j}-\tilde{b}_{j}\tilde{v}_{j}+\tilde{\alpha}_{j}\tilde{a}_{j})\end{cases},$ (13)
@@
where $\bar{g}_{i},\tilde{g}_{i}$ are the coupling functions, and $\epsilon$ is the coupling strength, We have chosen the coupling through nodes as the simplest and, possibly, physically feasible way to couple Dirac-Bianconi driven oscillators. In fact, we consider each Dirac-Bianconi driven oscillator as an independent dynamical unit exhibiting complex (oscillatory) dynamics induced by the Dirac-Bianconi coupling, while the coupling between the oscillators is modeled with a classic diffusive coupling via the variables on the nodes (0-cochains). In principle, we could consider more elaborated higher-order couplings, however a pairwise weak ( $\penalty 10000\ 10^{-2}$) coupling allows us to focus on the dynamics of the Dirac-Bianconi driven oscillators. We assume the functions $\bar{g}_{i}$ and $\tilde{g}_{i}$ to be linear and diffusive. For the setting depicted in Fig. 4, where the coupled variables are $\bar{u}_{1}$ and $\tilde{u}_{1}$, the only non-zero coupling functions are $\bar{g}_{i}=\tilde{u}_{1}-\bar{u}_{1}=-\tilde{g}_{i}$. Of course, different coupling configurations could be considered, as well as different kinds of pairwise couplings, including nonlinear and non-diffusive ones. Again, we choose a simple setting in order to illustrate the phenomenon and the method.
@@
As discussed previously, the two DBFHN oscillators have slightly different frequencies, meaning that an incoherent behavior is obtained without coupling, i.e., $\epsilon=0$. If we now set $\epsilon=10^{-2}$ and we consider Eqs. (12) and (13) coupled via only one node (as depicted in Fig. 4), we achieve synchronization only for large values of the coupling strength, as shown in Fig. 5, where we observe that the two oscillators synchronize their frequencies as the coupling becomes stronger. An analogous behavior is observed for other linear diffusive coupling combinations of different pairs of nodes (results not shown). As depicted in Fig. 5, the system undergoes a saddle-node bifurcation Crawford (1991); Strogatz (2018) with the coupling strength $\epsilon$ as the bifurcation parameter. In fact, when $\epsilon$ is large, there are at least two equilibrium points in the reduced phase model: one stable (a node), i.e., the synchronous state, and one unstable (a saddle). As the coupling strength diminishes, the two equilibrium points collide and annihilate each other. A more precise description of the unstable equilibrium point can be given via a phase description, as we will do in the next Section.
@@
Figure 5: Synchronization of two Dirac-Bianconi driven oscillators with weak diffusive coupling: we show the frequencies of DB1 (blue) and DB2 (red), coupled as in Eqs. (12) and (13) integrated until $t=6000$ t.u. and averaged over $100$ periods. We can observe that, with diffusive coupling, synchronization is achieved only for large values of the coupling strength. The parameters are $\bar{a}_{1}=\bar{a}_{3}=\tilde{a}_{1}=\tilde{a}_{3}=0.7$, $\bar{a}_{2}=\tilde{a}_{2}=-3$, $\bar{b}_{1}=\bar{b}_{2}=\bar{b}_{3}=\tilde{b}_{1}=\tilde{b}_{2}=\tilde{b}_{3}=0.3$, $\bar{\delta}_{1}=\bar{\delta}_{2}=\bar{\delta}_{3}=\tilde{\delta}_{1}=\tilde{\delta}_{2}=\tilde{\delta}_{3}=0.08$, $\bar{I}_{1}=\bar{I}_{2}=\bar{I}_{3}=\tilde{I}_{1}=\tilde{I}_{2}=0.8$ and $\tilde{I}_{3}=0.4$.
@@
In addition to the diffusive coupling, let us now consider also a Dirac-Bianconi coupling. We can impose that the diffusive-like coupling between 0-cochains has an effect on the $1$-cochains lying on the links (i.e., $1$-chains) adjacent to the coupled nodes (i.e., 0-chains); vice versa, the $1$-cochains have an effect on the 0-cochains lying on the nodes. If we look again at Fig. 4, this means that, via a Dirac-Bianconi coupling, $\bar{u}_{1}$ affects $\tilde{v}_{1}$ and $\tilde{v}_{3}$, and $\tilde{u}_{1}$ affects $\bar{v}_{1}$ and $\bar{v}_{3}$; vice versa $\tilde{v}_{1}$ and $\tilde{v}_{3}$ affect $\bar{u}_{1}$, and $\bar{v}_{1}$ and $\bar{v}_{3}$ affect $\tilde{u}_{1}$. We represent the higher-order coupling functions with $h^{\mathcal{D}}$.
@@
$\begin{cases}\displaystyle\dot{\bar{u}}_{i}=\bar{u}_{i}-\bar{u}_{i}^{3}-(\bar{\bm{B}}_{1}\vec{\bar{v}})_{i}+\bar{I}_{i}+\epsilon\bar{g}_{i}(\bar{u}_{i},\tilde{u}_{i_{1}},...,\tilde{u}_{i_{N_{n}}})+\epsilon\bar{h}^{\mathcal{D}}_{i}\Big(\bar{u}_{i},(\bar{\bm{B}}_{1}\vec{\tilde{v}})_{i_{1}},...,(\bar{\bm{B}}_{1}\vec{\tilde{v}})_{i_{N_{n}}}\Big),\\ \\ \displaystyle\dot{\bar{v}}_{j}=\bar{\delta}_{j}((\bar{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j}-\bar{b}_{j}\bar{v}_{j}+\bar{\alpha}_{j}\bar{a}_{j})+\epsilon\bar{h}^{\mathcal{D}}_{j}\Big(\bar{v}_{j},(\bar{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j_{1}},...,(\bar{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j_{N_{l}}}\Big),\end{cases}$ (14)
@@
$\begin{cases}\displaystyle\dot{\tilde{u}}_{i}=\tilde{u}_{i}-\tilde{u}_{i}^{3}-(\tilde{\bm{B}}_{1}\vec{\tilde{v}})_{i}+\tilde{I}_{i}+\epsilon\tilde{g}_{i}(\tilde{u}_{i},\bar{u}_{i_{1}},...,\bar{u}_{i_{N_{n}}})+\epsilon\tilde{h}^{\mathcal{D}}_{i}\Big(\tilde{u}_{i},(\tilde{\bm{B}}_{1}\vec{\bar{v}})_{i_{1}},...,(\tilde{\bm{B}}_{1}\vec{\bar{v}})_{i_{N_{n}}}\Big),\\ \\ \displaystyle\dot{\tilde{v}}_{j}=\tilde{\delta}_{j}((\tilde{\bm{B}}_{1}^{\top}\vec{\tilde{u}})_{j}-\tilde{b}_{j}\tilde{v}_{j}+\tilde{\alpha}_{j}\tilde{a}_{j})+\epsilon\tilde{h}^{\mathcal{D}}_{j}\Big(\tilde{v}_{j},(\tilde{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j_{1}},...(\tilde{\bm{B}}_{1}^{\top}\vec{\bar{u}})_{j_{N_{l}}}\Big).\end{cases}$ (15)
@@
In this case, the system achieves synchronization even for weak coupling strengths and, moreover, as the coupling strength grows, the oscillations are slowed down, as shown in Fig. 6. Observe that, qualitatively, the bifurcation behavior is the same as that of Fig. 5, but for a weaker coupling strength. Hence, in our setting the Dirac-Bianconi coupling [^6] Let us observe that the Dirac-Bianconi coupling of the previous equation would also emerge in the phase reduced model (see next Section) if the second order phase reduction León and Pazó (2019) of Eqs. (12) and (13) is performed, but it would be of order $\epsilon^{2}$. is fundamental in achieving synchronization for weak couplings. In the next section we will explain this behavior through the theory of phase reduction.
@@
Figure 6: Synchronization of two Dirac-Bianconi driven oscillators weakly coupled via the Dirac-Bianconi operator: we show the frequencies of DB1 (blue) and DB2 (red), coupled as in Eqs. (14) and (15) integrated until $t=6000$ t.u. and averaged over $100$ periods. On the left panel, we can observe that synchronization is achieved for small values of the coupling. The right panel shows that the Dirac-Bianconi coupling slows down the oscillations as the coupling strength is increased. Note the different scales of the $x$ axis in the two panels. The parameters are $\bar{a}_{1}=\bar{a}_{3}=\tilde{a}_{1}=\tilde{a}_{3}=0.7$, $\bar{a}_{2}=\tilde{a}_{2}=-3$, $\bar{b}_{1}=\bar{b}_{2}=\bar{b}_{3}=\tilde{b}_{1}=\tilde{b}_{2}=\tilde{b}_{3}=0.3$, $\bar{\delta}_{1}=\bar{\delta}_{2}=\bar{\delta}_{3}=\tilde{\delta}_{1}=\tilde{\delta}_{2}=\tilde{\delta}_{3}=0.08$, $\bar{I}_{1}=\bar{I}_{2}=\bar{I}_{3}=\tilde{I}_{1}=\tilde{I}_{2}=0.8$ and $\tilde{I}_{3}=0.4$.